OpenStudy (anonymous):

A shell from the uss missouri is fired at a 17 degree angle from the ground and has a range of 3.9km. find the original velocity of the shell in m/s

OpenStudy (irishboy123):

use this for range R \(\large R = \frac{v^2 \sin 2 \theta} {g} \) some good background to formula is here... https://en.wikipedia.org/wiki/Range_of_a_projectile ask if in doubt :p

OpenStudy (irishboy123):